OpenStudy (anonymous):

determine the value of 10x^-3 when x=1/2

OpenStudy (anonymous):

10x^-3=10/(x^3) = 10/((1/2)^3)=10/(1/8)=80

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I was assuming you did not leave off the parentheses like (10x)^-3

OpenStudy (anonymous):

no there are no parentheses

OpenStudy (anonymous):

i got 22 because 10 . 1/2 = 5 s^ = 25-3= 22

OpenStudy (anonymous):

how did u get 25

OpenStudy (anonymous):

\[10\times x^{-3}\] where x=1/2 is the question Supergirl exponents before multiplication